In the following situations, give the expected count for each of the k categories. (a) k = 3, H0: p1 = p2 = p3 = 1 3 , and n = 900

E1 =

E2 =

E3 =
Answer #1

Given,

If p1 = p2 = p3 = 13 , and n = 900

Expected count E(X) = n*p

E1 = p1*n = 11,700‬

E2 = p2*n = 11,700‬

E3 = p3*n = 11,700‬

If p1 = p2 = p3 = 1/3 and n = 900

E1 = p1*n = 300‬

E2 = p2*n = 300‬

E3 = p3*n = 300‬